OpenStudy (anonymous):

please help me..... The curve whose equation is y=ax^3 + bx^2 + cx + d has a point of inflexion at (-1,4), a turning point when x=2, and it passes through the point (3,-7). Find the values of a, b, c and d.

OpenStudy (anonymous):

a=3/11 b=9/11 c=-72/11 d=-23/11 \[y'=3ax ^{2}+2bx+c\] \[y''=6ax+2b\] inflection poit at (-1,4) means the second derivative =0 at x=1 which gives \[0=-6a+2b\] which means b=3a a turning point at x=2 means the first derivative =0 at x=2 which gives\[12a+4b+c=0\] So 24a+c=0 and c=-24a use the inf that f(3)=-7 and f(-1)=4 to solve for a and d